μ ≡ \frac{mass}{length}

velocity v = \sqrt{\frac{T}{μ}}

velocity v = \frac{ω}{k} = fλ

y_{1}(x, t) = A sin(kx − ωt)

y_{2}(x, t) = A sin(kx + ωt)

y_{1} + y_{2} = 2A sin(kx) cos(ωt)

2A sin(kx) = 0

nodes at x = 0, \frac{π}{k} , \frac{2π}{k} , \frac{3π}{k} , …, \frac{Nπ}{k}

nodes at x = 0, \frac{1}{2}λ, λ, \frac{3}{2}λ, …, \frac{N}{2} λ

2A sin(kx) = ± 2A

antinodes at x = \frac{π}{2k} , \frac{3π}{2k} , \frac{5π}{2k} , …, \frac{(2N + 1)π}{2k}

antinodes at x = \frac{1}{4}λ, \frac{3}{4}λ, \frac{5}{4}λ, …, \frac{(2N + 1)}{4}λ

μ_{s} =  ±  kg/m

v =  ±  m/s